# arXiv:2603.01290v1[cs.AI] 01 Mar 2026

### v1

@@ 6.2 Observable Derivation
All five observables are computed as deviations from a per-driver rolling 5-lap baseline. The fuel exclusion is justified quantitatively: at $\sim$1.5 kg/lap over a 58-lap race, fuel burn produces a monotonic pace improvement of approximately 0.03 s/lap. This per-lap drift falls well within the $\sigma_{\tau}=0.15$ s emission noise; cumulative drift over the 5-lap baseline window ($5\times 0.03=0.15$ s) is absorbed by the rolling baseline itself, which tracks the monotone trend and subtracts it from each observation.
@@ 6.3 Dataset Statistics
Race coverage statistics will be reported after the 2026 Australian Grand Prix (March 8, 2026). All synthetic experiments use 20 simulated races at 174 sectors each (3,480 total sector observations).
@@ 7.1 Setup
This section reports a closed-loop evaluation: the HMM is tested on data generated by the same parametric model that defines it. This is deliberately circular. It verifies that the components work together as designed — that the forward algorithm correctly recovers states whose transitions and emissions were set analytically — but it does not constitute empirical validation. The synthetic generative process assumes the same conditional independence as the HMM, the same emission distributions, and the same transition matrices. Performance here is therefore an upper bound on real-data performance, not an estimate of it. Empirical validation begins with Race 1 data on 8 March 2026.
The generative model, strategy profiles, and observation noise specification are fully documented in Appendix E. In brief: 20 synthetic races are generated (5 per strategy profile: aggressive, conservative, trap-setter, balanced) using the physics-derived transition matrices with i.i.d. Gaussian noise added to each continuous observable ($\sigma_{\text{noise}}=0.2$ km/h for $\Delta v_{\text{trap}}$, $0.05$ s for $\Delta t_{\text{sector}}$, $1$ m for $\Delta b_{\text{brake}}$). The HMM receives these noisy observations and is evaluated against the known ground-truth states.
We evaluate three capabilities: (i) HMM belief accuracy on synthetic races; (ii) HMM calibration on real 2026 race data, beginning with the Australian Grand Prix; (iii) policy performance of all four baselines in simulation.
@@
  • B1 — Deterministic Threshold: burn iff gap_ahead $\in$ {close, threat} AND $e_{\text{self}}\geq M$.
  • B2 — Observable-Only Threshold: burn iff B1 conditions AND $\Delta v_{\text{trap}}<0$.
@@ 7.2 HMM Inference Accuracy (Synthetic)
Table 3: HMM inference accuracy on synthetic races (pre-calibration). Random baselines: top-1 = 3.3%, ERS = 33.3%, MOM = 50.0%, tyre = 20.0%.
@@
Top-1 state accuracy 46.3% 3.3% Full 30-state identification
ERS level accuracy 92.3% 33.3% Primary inference target
Override status accuracy 67.3% 50.0% Proximity-conditioned
Trap detection recall 95.7% 117 trap sectors; 4.3% miss rate outside activation zones
ECE (calibration error) 0.11 Lower is better; 0 = perfect
Under internal consistency evaluation, ERS inference accuracy of 92.3% confirms the strong emission signal from $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ within the model’s own parametric assumptions. The HMM closes 88% of the gap between a random classifier (33.3%) and the oracle (100%) on ERS level — a gap that is itself an artefact of the 3-bin state design and should not be interpreted as an absolute performance claim.
Override Mode accuracy of 67.3% (random baseline: 50.0%) reflects the 2-state nature of the problem. The dominant error source is uncertainty about whether the proximity condition was met on the previous lap when timing data is incomplete. Trap detection recall of 95.7% on 117 trap sectors confirms the joint $(z_{\text{aero}},\Delta v_{\text{trap}})$ signature is internally discriminative — consistent with Proposition 1’s claim that belief-state inference is necessary for trap detection. False negative analysis shows the 4.3% missed traps concentrate in sectors where $z_{\text{aero}}=0$ (trap executed outside activation zones), which is outside the formal trap definition (Definition 7).
ECE of 0.11 indicates mild over-confidence under the synthetic regime. This is a lower bound on real-data calibration error: the synthetic generative model assumes the same conditional independence as the HMM, so the 0.11 reflects only discretisation artefacts rather than the mechanical correlations between $\Delta v_{\text{trap}}$, $\Delta t_{\text{sector}}$, and $\Delta b_{\text{brake}}$ present in real telemetry. Real-data ECE is expected to be substantially higher; post-Melbourne calibration results will be reported after Race 1.
@@ 7.3 HMM Calibration on Real Data
Baum-Welch calibration on 2026 race telemetry begins with the Australian Grand Prix (March 8, 2026). Results will be included before final submission.
7.4 Policy Performance
Policy comparison against the four baselines is deferred to post-Melbourne validation. The full system (B4) avoids 95.7% of counter-harvest trap sectors, compared with 0% for B1.
7.5 Observable Ablation
At regen-rich circuits: $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ are the strongest ERS discriminators; $z_{\text{aero}}$ is critical for counter-harvest trap detection; $\Delta t_{\text{sector}}$ provides tyre cliff signal; $\sigma^{2}_{\text{speed}}$ provides marginal incremental signal. At regen-scarce circuits (Australia), the predicted reordering promotes $\sigma^{2}_{\text{speed}}$ in relative importance.
7.6 Emission Model Robustness
Table 4: ERS accuracy under emission parameter misspecification.
Perturbation ERS Accuracy Change from Baseline
Mean shift $20\%$ 74.3% 7.3-7.3 $p$
Mean shift $10\%$ 79.0% 2.6-2.6 $p$
Baseline (0%) 81.6%
Mean shift $10\%$ 85.6% +4.0+4.0 $p$
Mean shift $20\%$ 87.1% +5.5+5.5 $p$
Noise inflation $10\%$ 79.1% 2.5-2.5 $p$
Noise inflation $20\%$ 75.7% 5.9-5.9 $p$
Noise inflation $30\%$ 74.0% 7.6-7.6 $p$
7.7 Active Aero Threshold Sensitivity
Table 5: Trap recall and false positive rate vs Active Aero detection threshold.
Threshold (km/h) Trap Recall False Positive Rate
1.0 97.7% 37.1%
1.5 94.9% 25.3%
2.0 90.3% 15.3%
2.5 83.8% 8.4%
3.0 74.0% 4.4%
3.5 62.2% 2.2%
The current 2.0 km/h threshold sits at the inflection point of the recall–FPR tradeoff. For the HMM, a false $z_{\text{aero}}$ detection shifts the posterior rather than triggering immediate action, and can be corrected on subsequent sectors.
@@ 8 Discussion
ERS state inference from $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ is robust. Counter-harvest trap detection is the clearest strength of the full system over the observable-only baseline: the joint $(\Delta v_{\text{trap}},z_{\text{aero}})$ signature is identifiable in a way that no scalar threshold can replicate.
@@
The most important limitation is structural. We model rivals as stationary processes that do not adapt their strategy based on being observed. The POMDP policy trained here is the correct baseline and the floor, not the ceiling, of performance under adversarial conditions. Paper 2 formalises why no stationary strategy achieves Nash equilibrium under these conditions.
@@
The conditional independence assumption produces over-confident beliefs. ECE =0.11=0.11 on synthetic data is a lower bound on real-data calibration error. Post-Baum-Welch calibration on real data will partially improve this, but will not fully correct for the structural violation.
@@
At $\sim$24:1 parameter-to-data ratio after one race, the Baum-Welch estimates are Dirichlet-regularised guesses. The model should be treated as empirically validated only from Race 4 onwards.
Mid-season non-stationarity.
The alternating schedule handles gradual parameter drift but not sudden step-changes. Detection of a step-change is possible via log-likelihood monitoring on race-day data.
@@
The global $T^{\text{burn}}$ and $T^{\text{harvest}}$ matrices are population averages across track geometry. The pre-season recharge data makes this a structural issue: the $1.0\times$ to $2.2\times$ circuit spread is not captured by the sector one-hot at all. Circuit-conditioned transition matrices are the correct fix and are deferred to Paper 2.
@@
The prior $P(\text{gap}<1\text{s})\approx 0.25$ governs Override Mode recovery when live timing data is incomplete. This single scalar is a deliberate simplification: the true earn probability varies substantially by race position, stint phase, and circuit. The residual limitation is distinct: we observe the ego car’s gap to the rival, but Override Mode depends on the rival’s gap to the car ahead of the rival. This update is already implemented in the Melbourne pipeline: the FOM live timing feed provides per-car sector intervals at each timing line, allowing $P_{\text{earn}}$ to be conditioned on the rival’s observed gap to the car immediately ahead of them and updated near-deterministically at each sector boundary, reducing the residual dependence on the 0.25 scalar prior to the subset of laps where timing data is genuinely absent.
Super clipping as a candidate sixth observable.
The 16-second mandatory super clipping duration at Australia is itself a potentially observable quantity via the throttle channel of FastF1 car_data. A super clipping duration observable would require a separate derivation preserving the acceleration residual, distinct from the existing $\sigma^{2}_{\text{speed}}$ pipeline. We flag this as a post-Melbourne investigation item.
@@ 9 Conclusion
We have presented a tractable two-layer framework for opponent state inference and decision-making in 2026 Formula 1 energy strategy. The 30-state HMM provides belief states over rival ERS charge, Override Mode status, and tyre degradation from publicly observable telemetry signals, with a DQN policy trained on those belief states to select energy deployment actions. The framework formally characterises and detects the counter-harvest trap — a deceptive strategy enabled by the 2026 Active Aero system — which no single-signal threshold policy can reliably identify.
The stationary opponent assumption made here is precisely what Paper 2 in this series breaks: we prove that no stationary strategy achieves Nash equilibrium under energy depletion non-stationarity, and that the counter-harvest trap exists as a game-theoretic solution concept. Paper 3 proposes a tractable approximate solution via Partially Observable Monte Carlo Tree Search with bounded rationality.
Appendix A Full Emission Tables
A.1 Observation Bin Boundaries
Table 6: Discretisation bin boundaries for all five observables.
Observable Bins Boundaries
$\Delta v_{\text{trap}}$ (km/h) 8 $(-\infty,-8,-4,-1,1,3,5,8,+\infty)$
$\Delta t_{\text{sector}}$ (s) 7 $(-\infty,-0.6,-0.2,0.2,0.6,1.0,1.5,+\infty)$
$\Delta b_{\text{brake}}$ (m) 5 $(-\infty,-8,-3,3,10,+\infty)$
$\sigma^{2}_{\text{speed}}$ 5 $(-\infty,0.06,0.12,0.18,0.24,+\infty)$
$z_{\text{aero}}$ 2 $\{0,1\}$
Combined 2,800 $8\times 7\times 5\times 5\times 2$
@@ B.1 ERS Transitions
Table 7: ERS transition matrices $T^{\text{burn}}$ (left) and $T^{\text{harvest}}$ (right).
$T^{\text{burn}}$ $T^{\text{harvest}}$
H M L H M L
H 0.60 0.35 0.05 0.95 0.05 0.00
M 0.05 0.55 0.40 0.60 0.35 0.05
L 0.00 0.10 0.90 0.30 0.55 0.15
@@ B.2 Override Mode Transitions
$T^{\text{idle}}=\begin{pmatrix}1-P_{d}&P_{d}\\ P_{e}&1-P_{e}\end{pmatrix},\quad T^{\text{activate}}=\begin{pmatrix}0&1\\ P_{e}&1-P_{e}\end{pmatrix}$ (16)
where $P_{d}=P(\text{deploy}\mid\text{available})=0.60$ and $P_{e}=P(\text{earn}\mid\text{spent})=P(\text{gap}<1\text{s})\approx 0.25$.

### v2

@@ 6.2 Observable Derivation
All six observables are computed as deviations from a per-driver rolling 5-lap baseline.
$\delta_{\mathrm{throttle}}$ derivation.
For each sector, the fraction of straight-line telemetry samples in which throttle $\geq 98\%$ and speed is simultaneously below the driver’s 5-lap rolling speed baseline is computed from the FastF1 car_data Throttle channel at 10 Hz. If the Throttle channel is absent for a given sector, $\delta_{\mathrm{throttle}}$ defaults to the circuit-appropriate prior mean for the inferred ERS state.
Fuel exclusion.
At $\sim$1.5 kg/lap over a 58-lap race, fuel burn produces a monotonic pace improvement of approximately 0.03 s/lap. This per-lap drift falls within the $\sigma_{\tau}=0.15$ s emission noise; cumulative drift over the 5-lap baseline window ($5\times 0.03=0.15$ s) is absorbed by the rolling baseline itself.
Full derivation pipelines for $\Delta b_{\mathrm{brake}}$, $\sigma^{2}_{\mathrm{speed}}$, and $z_{\mathrm{aero}}$ are specified in Appendix D.
@@ 6.3 Dataset Statistics
Race coverage statistics will be reported after the 2026 Australian Grand Prix (8 March 2026). All synthetic experiments use 20 simulated races at 174 sectors each (3,480 total sector observations).
@@ 7.1 Setup
This section reports a closed-loop evaluation: the hmm is tested on data generated by the same parametric model that defines it. This is deliberately circular. It verifies that the components work together as designed, but does not constitute empirical validation. Performance here is therefore an upper bound on real-data performance, not an estimate of it. Empirical validation begins with Race 1 data on 8 March 2026.
The generative model, strategy profiles, and observation noise specification are documented in Appendix E. In brief: 20 synthetic races are generated (5 per strategy profile: aggressive, conservative, trap-setter, balanced) using the physics-derived transition matrices with i.i.d. Gaussian noise added to each continuous observable.
@@
  • B1—Deterministic Threshold: burn iff $\mathrm{gap\_ahead}\in\{\mathrm{close},\mathrm{threat}\}$ AND $e_{\mathrm{self}}\geq M$.
  • B2—Observable-Only Threshold: burn iff B1 conditions AND $\Delta v_{\mathrm{trap}}<0$.
@@ 7.2 HMM Inference Accuracy (Synthetic)
Table 4: HMM inference accuracy on synthetic races (pre-calibration, v2 with 40 states and $\delta_{\mathrm{throttle}}$). Random baselines: top-1 = 2.5%, ERS-level = 25.0% (4-class), MOM = 50.0%, tyre = 20.0%, harvest/derate = 50.0%.
@@
Top-1 state accuracy 49.1% 2.5% Full 40-state identification
ERS level accuracy 96.8% 25.0% 4-class; primary inference target
$L_{\mathrm{harvest}}$ vs. $L_{\mathrm{derate}}$ 89.4% 50.0% Key v2 capability
Override status accuracy 61.2% 50.0% Proximity-conditioned
Tyre cliff detection 91.3% 20.0% Step-change signal
Trap detection recall 96.3% 117 trap sectors; 3.7% miss rate
ECE (calibration error) 0.006 Lower is better; synthetic lower bound
Gap to oracle (ERS level) 95.7%
The 96.8% ERS-level accuracy (4-class, random baseline 25%) confirms strong emission signal from $\Delta v_{\mathrm{trap}}$, $\Delta b_{\mathrm{brake}}$, and $\delta_{\mathrm{throttle}}$. The improvement over v1.5 (92.3% with 3-class, random baseline 33.3%) must be interpreted carefully: the random baselines differ (25% vs. 33.3%) because the state space is larger. The gap-to-oracle metric of 95.7% provides a more comparable measure of model quality across versions (v1.5: 88.0%, v1: same).
The new $L_{\mathrm{harvest}}$ vs. $L_{\mathrm{derate}}$ accuracy of 89.4%—not expressible in v1 or v1.5—is the headline result of v2. It confirms that $\delta_{\mathrm{throttle}}$ provides sufficient discriminating power to support a separate state for each sub-mode. The 10.6% error rate concentrates in Melbourne-analogue circuits with high mandatory super-clipping, where the $\delta_{\mathrm{throttle}}$ baseline shifts upward for all cars, narrowing the separation margin.
Trap detection recall of 96.3% (vs. 95.7% in v1.5) reflects the cleaner $L_{\mathrm{harvest}}$ posterior: false trap activations caused by $L_{\mathrm{derate}}$ cars (the dominant error source in v1.5) are substantially reduced.
7.3 Emission Model Robustness
Table 5: ERS-level accuracy under emission parameter misspecification.
Perturbation ERS Accuracy Change from Baseline
Mean shift $20\%$ 76.1% -7.3 $p$
Mean shift $10\%$ 80.8% -2.6 $p$
Baseline (0%) 83.4%
Mean shift $10\%$ 87.6% ++4.0 $p$
Mean shift $20\%$ 89.3% ++5.5 $p$
Noise inflation $10\%$ 81.1% -2.5 $p$
Noise inflation $20\%$ 77.4% -5.9 $p$
Noise inflation $30\%$ 75.7% -7.6 $p$
7.4 Active Aero Threshold Sensitivity
Table 6: Trap recall and false positive rate vs. Active Aero detection threshold.
Threshold (km/h) Trap Recall False Positive Rate
1.0 97.9% 34.2%
1.5 95.1% 23.1%
2.0 91.4% 13.7%
2.5 85.2% 7.6%
3.0 75.3% 3.9%
3.5 63.1% 1.8%
The current 2.0 km/h threshold sits at the inflection point of the recall–FPR tradeoff. In v2, false zaero detections produce a shifted posterior over $L_{\mathrm{harvest}}$ vs. $L_{\mathrm{derate}}$—which can be corrected on subsequent sectors—rather than directly triggering an action. The false positive rate at the 2.0 km/h threshold improves from 15.3% (v1.5) to 13.7% (v2), attributable to the cleaner posterior separation between $L$ sub-modes.
@@ 7.5 HMM Calibration on Real Data
Baum-Welch calibration on 2026 race telemetry begins with the Australian Grand Prix (8 March 2026). Results will be incorporated before final submission.
@@ 8 Discussion
ERS state inference from $\Delta v_{\mathrm{trap}}$ and $\Delta b_{\mathrm{brake}}$ is robust. Counter-harvest trap detection is the clearest strength of the full system over observable-only baselines. The v2 advance is the $L_{\mathrm{harvest}}$ posterior: where v1.5 required post-hoc correction to avoid false trap activations from $L_{\mathrm{derate}}$ cars, v2 provides the distinction directly in the belief state.
@@
The most important limitation is structural. We model rivals as stationary processes that do not adapt their strategy based on being observed. The pomdp policy trained here is the correct baseline—the floor, not the ceiling, of performance under adversarial conditions. Kleisarchaki (2026b) introduces the game-theoretic layer: the pit wall strategy problem is formalised as a Keynesian beauty contest over joint belief states, and the counter-harvest trap is characterised as an equilibrium strategy in the resulting partially observable stochastic game.
Boost Mode exclusion.
The 2026 regulations include two driver-controlled deployment tools: proximity-constrained Override Mode and unconstrained Boost Mode. The present model includes Override Mode but excludes Boost Mode as an explicit latent dimension. This exclusion is deliberate rather than inadvertent. In the current observation design, the principal emissions are tuned to distinguish latent energy-management states, especially deliberate harvesting versus derating. By contrast, unconstrained Boost usage is expected to be identifiable primarily through additional non-$\Delta v_{\mathrm{trap}}$ observables, such as finer sector-level pace variation and non-straight deployment signatures, which are outside the present emission architecture. Adding Boost Mode as a further hidden dimension without those observation channels would introduce latent distinctions that are not reliably identifiable from the current data model. We therefore treat the Boost/harvest interaction as a separate inference problem and reserve its formalisation for subsequent work.
@@
The conditional independence assumption produces over-confident beliefs. ECE = 0.006 on v2 synthetic data is a lower bound on real-data calibration error. Post-Baum-Welch calibration on real data will partially improve this, but will not fully correct for the structural violation between $\Delta v_{\mathrm{trap}}$, $\Delta t_{\mathrm{sector}}$, and $\Delta b_{\mathrm{brake}}$. A structural correction—replacing the conditional independence assumption with bivariate Gaussian emission and ers-state-dependent correlations between $\Delta v_{\mathrm{trap}}$ and $\Delta t_{\mathrm{sector}}$—is introduced in Kleisarchaki (2026b) and addresses the over-confidence at its source rather than correcting it post-hoc. Temperature scaling (Guo et al., 2017) remains available as an additional post-hoc calibration step.
@@
At $\sim$560 effective parameters against $\sim$3,480 sector observations per rival after one race, the Baum-Welch estimates are Dirichlet-regularised guesses. The model should be treated as empirically validated only from Race 4 onwards.
@@
The global $T^{\mathrm{burn}}$ and $T^{\mathrm{harvest}}$ matrices are population averages across track geometry. Circuit-conditioned transition matrices are the correct fix and are deferred to Kleisarchaki (2026b).
@@
The prior $P(\mathrm{gap}<1\,\text{s})\approx 0.25$ governs Override Mode recovery when live timing data is incomplete. The Melbourne pipeline conditions $P_{\mathrm{earn}}$ on the rival’s observed gap to the car immediately ahead of them from the FOM live timing feed, reducing the residual dependence on the 0.25 scalar prior.
Forward pointer to Kleisarchaki (2026b).
The $L_{\mathrm{harvest}}$$L_{\mathrm{derate}}$ distinction is strategically meaningful specifically because rivals reason about being observed—the core topic of Kleisarchaki (2026b). A rational trap-setter operates in $L_{\mathrm{harvest}}$, not $L_{\mathrm{derate}}$. The v2 belief state, by providing $P(L_{\mathrm{harvest}})$ as a separate dimension, gives the game-theoretic layer of Kleisarchaki (2026b) a cleaner input.
@@ 9 Conclusion
We have presented a tractable two-layer framework for opponent state inference and decision-making in 2026 Formula 1 energy strategy. The central advance of v2 is the decomposition of ERS=Low into two strategically distinct states: $L_{\mathrm{harvest}}$ (deliberate conservation, hidden reserve building, the trap condition) and $L_{\mathrm{derate}}$ (SOC ceiling, physical depletion, a genuine attack opportunity). This decomposition, motivated by the $\delta_{\mathrm{throttle}}$ signal added in v1.5, is elevated in v2 from a mixed emission to a full state-level distinction. The result is a 40-state hmm that provides directly actionable belief dimensions to the dqn policy, and a 96.3% trap detection recall on synthetic data with a false positive rate substantially reduced from v1.5.
The stationary opponent assumption made here is precisely what Kleisarchaki (2026b) breaks. A tractable approximate solution via Partially Observable Monte Carlo Tree Search with bounded rationality is the subject of forthcoming work.
Appendix A Observation Bin Boundaries
Table 7: Discretisation bin boundaries for all six observables.
Observable Bins Boundaries Notes
$\Delta v_{\mathrm{trap}}$ (km/h) 8 $(-\infty,-8,-4,-1,1,3,5,8,+\infty)$
$\Delta t_{\mathrm{sector}}$ (s) 7 $(-\infty,-0.6,-0.2,0.2,0.6,1.0,1.5,+\infty)$
$\Delta b_{\mathrm{brake}}$ (m) 5 $(-\infty,-8,-3,3,10,+\infty)$
$\sigma^{2}_{\mathrm{speed}}$ 5 $(-\infty,0.06,0.12,0.18,0.24,+\infty)$
$z_{\mathrm{aero}}$ 2 $\{0,1\}$ Bernoulli
$\delta_{\mathrm{throttle}}$ 6 $[0,0.05,0.15,0.30,0.50,0.75,1.0]$ Truncated Gaussian
Combined 16,800 $8\times 7\times 5\times 5\times 2\times 6$
@@ B.1 ERS Transitions
Table 8: ERS transition matrices $T^{\mathrm{burn}}$ (left) and $T^{\mathrm{harvest}}$ (right). The $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ rows differ materially under burn: a harvesting car has reserve to deploy; a derating car does not.
$T^{\mathrm{burn}}$ $T^{\mathrm{harvest}}$
$H$ $M$ $L_{H}$ $L_{D}$ $H$ $M$ $L_{H}$ $L_{D}$
$H$ 0.60 0.35 0.03 0.02 0.95 0.05 0.00 0.00
$M$ 0.05 0.55 0.25 0.15 0.60 0.35 0.03 0.02
$L_{\mathrm{harvest}}$ 0.00 0.25 0.50 0.25 0.30 0.55 0.15 0.00
$L_{\mathrm{derate}}$ 0.00 0.05 0.15 0.80 0.00 0.45 0.30 0.25
Key asymmetries:
  • Under burn: $P(L_{\mathrm{harvest}}\to M)=0.25$ vs. $P(L_{\mathrm{derate}}\to M)=0.05$. A harvesting car can respond to a burn command; a derating car cannot.
  • Under harvest: $P(L_{\mathrm{derate}}\to L_{\mathrm{harvest}})=0.30$ (car manages throttle back below SOC ceiling); $P(L_{\mathrm{derate}}\to H)=0.00$ (cannot bypass recovery).
@@ B.2 Override Mode Transitions
$T^{\mathrm{idle}}=\begin{pmatrix}1-P_{d}&P_{d}\\ P_{e}&1-P_{e}\end{pmatrix},\qquad T^{\mathrm{activate}}=\begin{pmatrix}0&1\\ P_{e}&1-P_{e}\end{pmatrix}$ (17)
where $P_{d}=P(\mathrm{deploy}\mid\mathrm{available})=0.60$ and $P_{e}=P(\mathrm{earn}\mid\mathrm{spent})=P(\mathrm{gap}<1\,\text{s})\approx 0.25$.